Two firms produce the same product in a local market where the demand of product is P = 780−2Q - P is the price of the product and Q is the total amount of product produced by these two firms. Each firm needs to decide the quantity that they are going to produce given the local demand, and the cost of producing one unit is $1.

Write down the normal-form representation of the game.
Answer #1

It is given that there are 2 firms producing the same product. The demand function is given as P = 780 - 20. The MC of producThe profit of each firm is calculated below: Profit = (P-MC) Quantity = ($390.5 -1)×97.375 = $37,927.56 Thus, the profit earn779–292 = 2(29) 779–292 = 2(779-492) 779 – 292 = 1,558 - 892 -29, +8q, = 1,558 – 779 692 = 779 779 926 91 = 72 = 129.83 Thus,The normal-form representation of game is as follows: Q2 Production 97.375 129.83 97.375 37927.56, 37927.56 4 2141.5, 31606 1